# Multimodal Large Language Models to Support Real-World Fact-Checking

### v1

@@ Abstract
Multimodal large language models (MLLMs) carry the potential to support humans in processing vast amounts of information. While MLLMs are already being used as a fact-checking tool, their abilities and limitations in this regard are understudied. Here is aim to bridge this gap. In particular, we propose a framework for systematically assessing the capacity of current multimodal models to facilitate real-world fact-checking. Our methodology is evidence-free, leveraging only these models’ intrinsic knowledge and reasoning capabilities. By designing prompts that extract models’ predictions, explanations, and confidence levels, we delve into research questions concerning model accuracy, robustness, and reasons for failure. We empirically find that (1) GPT-4V exhibits superior performance in identifying malicious and misleading multimodal claims, with the ability to explain the unreasonable aspects and underlying motives, and (2) existing open-source models exhibit strong biases and are highly sensitive to the prompt. Our study offers insights into combating false multimodal information and building secure, trustworthy multimodal models. To the best of our knowledge, we are the first to evaluate MLLMs for real-world fact-checking.
@@ 1 Introduction
Misinformation is a significant challenge on the internet, especially with regard to multimodal claims, which combine text, images, videos, and other media types Mubashara et al. (2023). The visual component in these cases could be manipulated, used out-of-context, or accompanied by text that makes a false claim Huh et al. (2018); Luo et al. (2021); Aneja et al. (2023); Yao et al. (2023). In such cases, fact-checkers and the tools they employ need to be able to handle multiple modalities.
@@
Large language models (LLMs) store extensive information about the world, which exceeds what any individual human can know and is accessible in a far more human-friendly way than any search engine can offer Zhao et al. (2023); Hu et al. (2023a). As such, they can be a powerful tool in the hands of fact-checkers, who often seek extra factual knowledge to verify a claim Nakov et al. (2021). The application of LLMs to fact-checking has been extensively studied in a text-only setting, wherein a model receives as input both a textual claim, with or without a set of evidence documents, and needs to reason to determine the veracity of the claim Chen et al. (2022); Pan et al. (2023); Li et al. (2023a). LLMs have increasingly strong understanding and reasoning capabilities, so using them to sift through evidence documents when verifying a claim is proving very effective Chen et al. (2022); Li et al. (2023a). Yet, this pipeline approach requires an extra step of evidence retrieval, which is computationally heavy and error-prone, while it overlooks the vast amount of knowledge stored in the parameters of LLMs Hu et al. (2023b). Additionally, LLMs are now giving way from text-only to multimodal models, which can process signals from multiple sources and perform a wider range of tasks Liu et al. (2023b); Zhang et al. (2023); Yin et al. (2023). We hypothesize that multimodal large language models (MLLMs) trained on large amounts of data can serve as a sufficient substitute for retrieving evidence and perform fact-checking in an evidence-free fashion, i.e., relying solely on their parametric knowledge.
@@
Despite their great promise, however, MLLMs have limitations as well, specifically with respect to factuality Fu et al. (2023); Liu et al. (2023c); Xu et al. (2023). Models may produce hallucinations Huang et al. (2023); Li et al. (2023b) and suffer from a lack of knowledge or exhibit biases Hu et al. (2023a); Wang et al. (2023). To understand MLLMs’ potential to support fact-checking real-world claims, we propose the evaluation framework illustrated in Figure 1. It involves querying models for predictions, explanations, and confidence scores on multimodal check-worthy claims. Based on the data obtained, we can further assess the models’ capabilities from various perspectives, addressing the following research questions:
@@
RQ1: How accurate are MLLMs in identifying misinformation in multimodal claims?
RQ2: How do different modalities (text and image) contribute to the assessment of multimodal content?
RQ3: Are models sensitive to prompts in different languages or expressions, or do they exhibit robustness?
RQ4: Can in-context learning enhance model performance?
RQ5: Can MLLMs accurately estimate the correctness of their responses in terms of a confidence score?
@@
RQ6: What types of errors are most common for MLLMs employed as multimodal fact-checkers?
Our investigation spans an array of vision-language models, including GPT-4V, LLaVA-1.5 Liu et al. (2023b), MiniGPT-v2 Chen et al. (2023), and InstructBLIP Dai et al. (2023), which stand as leaders in this domain. We evaluate these models across three existing multimodal fact-checking datasets Zlatkova et al. (2019); Aneja et al. (2023); Yao et al. (2023), and one new dataset, which we build from recent check-worthy claims to mitigate the risk of data contamination.
Our study finds that MLLMs display a range of fact-checking capabilities. GPT-4V, in particular, excels in our tests, exhibiting high accuracy, useful explanations, and good calibration, with the accuracy on different datasets generally reaching around . On the other hand, open-source models like MiniGPT-v2 and InstructBLIP lack explanatory and uncertainty-reporting capabilities. LLaVA models show a strong sensitivity to prompt formulation, a tendency to comply with the question in the prompt, and hallucinatory behaviors.
@@
In summary, we systematically evaluate existing multimodal models to determine their capability to support real-world fact-checking. We formulate several basic research questions and assess their accuracy, robustness, calibration, and reasons for failure across various datasets. Our work fills a research gap, providing a comprehensive overview of the domain and shedding light on both fact-checking and MLLMs’ factuality. As far as we are aware, our study is the first assessment of MLLMs in the context of real-world fact-checking.
2 Background and Related Work
@@
Hoes et al. (2023) find that ChatGPT accurately classifies 69% of statements in a dataset built from the fact-checking platform PolitiFact. FactLLaMA Cheung and Lam (2023) integrates external evidence into the instruct-tuning process to enhance the model’s ability to leverage evidence for predictions. Studies have also focused on leveraging the capabilities of LLMs to develop automated frameworks for decomposing claims, generating queries, and predicting based on gathered evidence, thereby augmenting the interpretability of the fact-checking process Pan et al. (2023); Li et al. (2023a). However, these works often focused solely on prediction accuracy. Our work analyzes the model’s fact-checking capabilities from three perspectives: predictions, explanations, and confidence.
@@
Multimodal misinformation has been found to spread faster than its text-only counterparts Li and Xie (2020). It primarily falls into two categories: one involves textual claims about manipulated content across different modalities, while the other pertains to out-of-context misinformation, featuring unaltered images, audio, or videos in misleading contexts Luo et al. (2021); Aneja et al. (2023); Yao et al. (2023). The multimodal fact-checking is gnerally structured into distinct stages within pipelines, including claim detection extraction, evidence retrieval, verdict prediction, and explanation Nakov et al. (2021); Mubashara et al. (2023). In our work, we posit that evidence is parametrically encoded within the models, enabling us to bypass the evidence retrieval phase and directly engage in predictions and explanations for the targeted claims.
Evaluation of MLLMs
With the advancement of MLLMs, the systematic evaluation of their capabilities has become increasingly critical. A broad array of benchmarks has been established to estimate the MLLM’s performance across a variety of tasks, including perceptual tasks such as image recognition, image description, and optical character recognition, as well as cognitive tasks like commonsense reasoning and numerical calculation Yin et al. (2023); Fu et al. (2023); Liu et al. (2023c); Xu et al. (2023); Tong et al. (2024). Tu et al. (2023) access the resilience of MLLMs against various image-related attacks from a security perspective. However, there is an absence of research on evaluating the ability to conduct multimodal fact-checking with MLLMs. Our work aims to bridge this gap.
@@ 3 Evaluation Framework
We propose an evaluation framework that analyzes model capabilities along the various dimensions illustrated in Figure 1. Our experiments are conducted in an evidence-free setting, meaning that models have access only to the multimodal claims without any external knowledge. Through extensive controlled experiments, we conduct a detailed analysis of models’ potential in fact-checking real-world claims, including their accuracy and reliability.
@@ 3.1 Evaluation Prompt
We simultaneously obtain the predictions, explanations, and confidence levels from MLLMs with the prompt below. This prompt is formulated following the example of prompts found in related fact-checking work Pan et al. (2023); Min et al. (2023), while for acquiring explanations and confidence estimates we refer to Xiong et al. (2023).
@@
3.2 Metrics
@@
Based on the three components (prediction, explanation, and confidence score) in the model’s response, we categorize them into four types: True, False, Uncertain, and Deny. Deny refers to instances where the model fails to respond according to our requirements, missing any component. Uncertain denotes cases where the model explicitly expresses uncertainty, with statements such as “cannot confirm”, “cannot verify”, or “cannot determine”, etc. Others are classified as True or False based on the model’s stated prediction. This classification scheme is designed to reveal the model’s inherent biases through the distribution of True and False responses, ability to follow instructions the number of Deny responses, and assess its knowledge-awareness through the frequency of Uncertain cases.
@@
Assuming the total count of True (T), False (F), Uncertain (U) and Deny (D) responses is . Since not all responses include predictions, we denote the number of samples containing predictions as , with . We introduce two distinct accuracy metrics, True/False Accuracy is defined as , and Overall Accuracy is defined as , where represent the number of correctly predicted samples within the corresponding group. The first metric posits that predictions made with expressed uncertainty should be omitted from accuracy assessments. The second metric provides a holistic prediction-based score over the same set of data points for all models, thus allowing for a fair comparison.
3.3 Models

### v2

@@ Abstract
Misinformation poses a real-world threat, especially when combined with misleading images. Multimodal large language models (MLLMs), which combine image comprehension with the rich knowledge and explanatory capability of language models, have become tools for humans to process vast information. However, their capacity and limitations as multimodal tools in assisting with fact-checking remain understudied. Here is aim to bridge this gap. In particular, we propose a framework for systematically assessing the capacity of current multimodal models to facilitate real-world fact-checking. Our methodology is evidence-free, leveraging only these models’ intrinsic knowledge and reasoning capabilities. By designing prompts that extract models’ predictions, explanations, and confidence levels, we conduct a detailed analysis of the model’s accuracy, bias, and other key factors. We empirically find that (1) GPT-4V exhibits surprising performance across various datasets, with an accuracy rate exceeding , and is capable of providing impressive explanations, and (2) even with the aid of prompt ensembles and in-context learning, open-source models significantly lag in performance. However, they show potential in remembering checked claims and reasoning out manipulated images. We also summarize the failure reasons, which contributes to strategies for future improvements. Our study offers insights into leveraging MLLMs to combat multimodal misinformation.
@@ 1 Introduction
Misinformation is a significant challenge on the internet, especially with regard to multimodal claims, which combine text, images, videos, and other media types Akhtar et al. (2023). The visual component in these cases could be manipulated or used out-of-context (OOC) to make a false claim Huh et al. (2018); Luo et al. (2021); Aneja et al. (2023); Yao et al. (2023). In such cases, fact-checkers and the tools they employ need to be able to handle multiple modalities.
@@
Large language models (LLMs) store extensive information, which exceeds what any individual human can know and is accessible in a far more human-friendly way than any search engine can offer Zhao et al. (2023); Hu et al. (2024). As such, they can be a powerful tool in the hands of fact-checkers, who often seek extra factual knowledge to verify a claim Nakov et al. (2021). The application of LLMs to fact-checking has been extensively studied in a text-only setting, wherein a model receives as input both a textual claim, with or without a set of evidence documents, and needs to reason to determine the veracity of the claim Chen et al. (2022); Pan et al. (2023); Li et al. (2023a). LLMs have increasingly strong understanding and reasoning capabilities, so using them to sift through evidence documents when verifying a claim is proving very effective Chen et al. (2022); Li et al. (2023a). Yet, this pipeline approach requires an extra step of evidence retrieval, which is computationally heavy and error-prone, while it overlooks the vast amount of knowledge stored in the parameters of LLMs Hu et al. (2023). We hypothesize that multimodal large language models (MLLMs) trained on large amounts of data can serve as a sufficient substitute for the retrieval of evidence and perform fact-checking in an evidence-free fashion, i.e., relying solely on their parametric knowledge.
@@
Despite their great promise, MLLMs have limitations as well, specifically with respect to factuality Fu et al. (2023); Liu et al. (2023c). Models may produce hallucinations Huang et al. (2023); Li et al. (2023b) and suffer from a lack of knowledge or exhibit biases Hu et al. (2024); Wang et al. (2023). To understand MLLMs’ potential to support fact-checking real-world claims, we propose the evaluation framework illustrated in Figure 1. It involves querying models for predictions, explanations, and confidence scores on multimodal check-worthy claims. Based on the data obtained, we can further assess the models’ capabilities from various perspectives, addressing the following research questions:
@@
RQ1 Performance Evaluation: How good are MLLMs in identifying multimodal misinformation? Can they perform accurate reasoning?
RQ2 Improving Approaches: How can we effectively boost the model’s fact-checking capabilities? Can the improvements be reflected in the reasoning, beyond accuracy metrics?
@@
RQ3 Taxonomy of Failure Reasons: What are the typical errors for MLLMs employed as multimodal fact-checkers?
Our investigation spans an array of vision-language models, including GPT-4V, MiniGPT-v2 Chen et al. (2023) and LLaVA-1.5 Liu et al. (2023b), which stand as leaders in this domain. We evaluate these models across three existing multimodal fact-checking datasets: Fauxtography Zlatkova et al. (2019), COSMOS Aneja et al. (2023), MOCHEG Yao et al. (2023), and one new dataset, which build from recent check-worthy claims to mitigate the risk of data contamination. We use prompt ensembles (PE) and in-context learning (ICL) to help improve the open-source MLLMs. We do not use fine-tuning because models are required to explain its reasoning process and provide confidence scores, while high-quality annotated data is scarce. Our study finds that MLLMs display a range of fact-checking capabilities. GPT-4V, in particular, excels in our tests, exhibiting high accuracy, useful explanations, and good calibration. The accuracy on various datasets generally reaches around . MiniGPT-v2 lacks explanatory and uncertainty-reporting capabilities. LLaVAs can follow instructions, and both PE and ICL enhance their performance to varying degrees, with the latter providing greater improvement. On Fauxtography and COSMOS, ICL increases the Overall Accuracy from 52.3% to 62.3% and from 67.7% to 76.5%, respectively. The majority baseline for both datasets is 50%.
@@
In summary, we systematically evaluate existing multimodal models to determine their capability to support real-world fact-checking. Our approach is more comprehensive than prior studies, which are limited to specific aspects such as face spoofing, forgery detection, or out-of-context detection Shi et al. (2024); Qi et al. (2024). We formulate several research questions to evaluate these models based on accuracy, bias, reasoning capabilities, and errors across a variety of datasets. Our work fills a research gap, providing a comprehensive overview of the domain and shedding light on leveraging MLLMs to combat multimodal misinformation.
2 Related Work
@@
Caramancion (2023) compared the performance of models such as ChatGPT 3.5 and ChatGPT 4.0 in news fact-checking.FactLLaMA Cheung and Lam (2023) integrates external evidence into the instruct-tuning process to enhance the model’s ability to leverage evidence for predictions. Studies have also focused on leveraging the capabilities of LLMs to develop automated frameworks for decomposing claims, generating queries, and predicting based on gathered evidence, thereby augmenting the interpretability of the fact-checking process Pan et al. (2023); Li et al. (2023a). However, these works often focused solely on prediction accuracy. Our work analyzes the model’s fact-checking capabilities from three perspectives: predictions, explanations, and confidence.
@@
Multimodal misinformation primarily falls into two categories: one involves textual claims about manipulated content across different modalities, while the other pertains to out-of-context misinformation, featuring unaltered images, audio, or videos in misleading contexts Luo et al. (2021); Aneja et al. (2023); Yao et al. (2023). Shi et al. (2024) proposed a benchmark to evaluate the performance of MLLMs in detecting face spoofing and forgery. Qi et al. (2024) introduced Sniffer, an MLLM designed for OOC detection and explanation, utilizing fine-tuning to boost the accuracy in identifying celebrities and clarifying inconsistencies between text and images. In our work, we posit that evidence is parametrically encoded within the models, enabling us to bypass the evidence retrieval phase and directly engage in predictions and explanations for the real-world claims.
@@ 3 Evaluation Framework
We propose an evaluation framework that includes datasets, prompts, and evaluation metrics to address the research questions.
3.1 Datasets
Due to constraints posed by the GPT-4V API, such as decreased query speeds after exceeding a daily limit, we sample data from various datasets and preprocess them to fit our evaluation framework. Additionally, we develop a new dataset to ensure the model has not encountered these data.
Fauxtography
is a multimodal fact-checking dataset sourced from the websites Snopes and Reuters Zlatkova et al. (2019). We select a random subset of 400 entries, ensuring a balanced composition of 200 true and 200 false instances.
COSMOS
is a dataset built from Snopes and News Outlets, and tailored for OOC detection Aneja et al. (2023). We randomly select 240 true and 240 false samples from the test split. Each image in this dataset comes with two captions, and we manually select one caption per image to ensure it aligns with our criteria.
MOCHEG
Post-4V
We collect new data from Snopes, focusing on articles published after the release of GPT-4V (September 26, 2023). We apply the same filtering methodology used in MOCHEG, resulting in a dataset of 186 samples.
| Fauxtography | COSMOS | MOCHEG | Post-4V | |
| True | 200 | 240 | 267 | 81 |
| False | 200 | 240 | 237 | 105 |
| Total | 400 | 480 | 504 | 186 |
@@ 3.2 Evaluation Prompt
We simultaneously obtain the predictions, explanations, and confidence levels from MLLMs with the prompt below. This prompt is formulated following the example of prompts found in related fact-checking work Pan et al. (2023); Min et al. (2023), while for acquiring explanations and confidence estimates we refer to Xiong et al. (2024). We collect their verbalized confidence as it is increasingly used for decision-making during human-machine collaboration Geng et al. (2023).
{adjustwidth}0.5cm0.5cm
@@
| Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
| T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | |
| GPT-4V | 158 | 195 | 29 | 18 | 81.9 | 82.1 | 179 | 204 | 83 | 14 | 86.2 | 80.0 | 216 | 223 | 37 | 28 | 87.2 | 83.5 | 54 | 98 | 26 | 8 | 79.6 | 73.7 |
| MiniGPT | - | - | - | - | - | 55.5 | - | - | - | - | - | 62.1 | - | - | - | - | - | 54.3 | - | - | - | - | - | 55.9 |
| LLaVA(7b) | 337 | 6 | 1 | 56 | 53.4 | 46.3 | 449 | 7 | 4 | 20 | 52.0 | 50.0 | 409 | 6 | 0 | 89 | 54.2 | 44.8 | 157 | 1 | 3 | 25 | 42.0 | 37.1 |
| LLaVA(13b) | 286 | 98 | 3 | 13 | 54.4 | 52.3 | 331 | 134 | 13 | 2 | 69.7 | 67.7 | 363 | 112 | 28 | 1 | 61.5 | 58.1 | 116 | 65 | 2 | 3 | 56.4 | 55.9 |
| LLaVA+PE | 244 | 153 | 2 | 1 | 57.1 | 54.7 | 275 | 204 | 0 | 1 | 76.3 | 71.7 | 290 | 214 | 0 | 0 | 59.9 | 58.1 | 85 | 101 | 0 | 0 | 56.9 | 56.1 |
| LLaVA+ICL-1 | 228 | 159 | 6 | 8 | 61.8 | 62.3 | 293 | 175 | 9 | 3 | 74.8 | 74.1 | 254 | 240 | 7 | 4 | 62.4 | 62.5 | 79 | 101 | 5 | 1 | 58.7 | 57.5 |
| LLaVA+ICL-2 | 186 | 188 | 8 | 18 | 61.6 | 61.7 | 247 | 215 | 12 | 7 | 77.3 | 76.5 | 195 | 286 | 10 | 13 | 60.2 | 60.4 | 48 | 122 | 8 | 8 | 62.1 | 61.4 |
| Majority | - | - | - | - | - | 50.0 | - | - | - | - | 50.0 | - | - | - | - | - | 53.0 | - | - | - | - | - | 56.5 | |
3.3 Evaluation Metrics
@@
Based on the three components (prediction, explanation, and confidence level) in the model’s response, we categorize them into four types. Others refer to instances where the model fails to respond according to our requirements, missing any component. Uncertain denotes cases where the model explicitly expresses uncertainty, with statements such as “cannot confirm”, “cannot verify”, or “cannot determine”, etc. The remaining samples are grouped as True or False according to the model’s prediction. Overall, the numbers of these four indicators reveal the model’s ability to follow instructions, express uncertainty, and inherent bias in predicting True and False.
@@
We design two accuracy metrics to reflect the model’s performance. One is True&False Accuracy, measuring the accuracy of samples that are solely classified as True or False. This is important because users usually do not rely on the model’s response when the model clearly expresses uncertainty or does not follow instructions. The other metric, named Overall Accuracy, calculates the proportion of responses that contain correct predictions across all samples. This facilitates comparison between different approaches.
4 Experimental Setups